skippy leaves on an highland where she produces two goods. According to the production possibility frontier 100=x2+y2 and she consumes all goods herself. A utility function is U=xy. Find a utility maximizing X and Y as well as the value of lamda.


1
Expert's answer
2021-07-15T04:33:43-0400

Create a new equation form the original information


"L=xy-\lambda(x^2+y^2-100)"

"\dfrac{\partial L}{\partial x}=y-2\lambda x=0"

"\dfrac{\partial L}{\partial y}=x-2\lambda y=0"

"\dfrac{\partial L}{\partial \lambda}=-(x^2+y^2-100)=0"

"y=2\lambda x"

"x=4\lambda^2x"


"x^2+4\lambda^2x^2-100=0"

"x>0"


"y=2\lambda x"

"4\lambda^2=1"


"x^2+x^2-100=0"

"x\not=0"



If "\lambda=-\dfrac{1}{2}," then "x" and "y" have opposite signs. So "U=xy<0"

If "\lambda=\dfrac{1}{2}," then


"x=5\sqrt{2}, y=\dfrac{5\sqrt{2}}{2}"

"x=5\sqrt{2}, y=\dfrac{5\sqrt{2}}{2}, \lambda=\dfrac{1}{2}"


"U_{max}=25"